OpenStudy (deadshot):

What is the secant of the angle in standard position in which the terminal side passes through the point (-16, -30)?

OpenStudy (deadshot):

I know that the angle is approximately 28°

OpenStudy (phi):

How did you get 28 degrees? Here is a plot

OpenStudy (phi):

the way to do these problems is 1) plot the point 2) make a triangle. the vertical side of the triangle will have length= y value= -30 the horizontal side of the triangle will have length= x value = -16 3) find the hypotenuse using c^2 = a^2 + b^2. here c^2= 900+256= 1156, and c= sqr(1156)= 34 4) now use SOH CAH TOA to find the sin, cos or tan as required 5) if you need secant, use secant= 1/cos and cosecant= 1/sin

OpenStudy (deadshot):

thanks!